OpenStudy (anonymous):

The length of the altitude drawn to the hypotenuse of a right triangle is 8. What could be the possible lengths of the segments of the hypotenuse? 4 and 4 4 and 8 4 and 16 4 and 64

OpenStudy (y2o2):

i think all answers are wrong

OpenStudy (anonymous):

Those were the answer choices given to me..

OpenStudy (anonymous):

THM: alt to hyp, alt is geomean b/w sgmts of hyp. So, 4/8 = 8/4 nope 4/8 = 8/8 nope 4/8 = 8/16 YES! 4/8 = 8/64 nope Answer is 4 and 16.

OpenStudy (anonymous):

thanks :)